OpenStudy (anonymous):

I am so confused! Question is: Write each expression in simplest form. Examples that get me lost y^ (5/8) divide y^(1/2) x^(1/4)*x^1/6*x^(1/3) (m^3n^-5*m^-1)^-3

OpenStudy (anonymous):

I don't even know what to begin and start with

OpenStudy (unklerhaukus):

\[a^n\times a^m=a^{n+m}\] \[a^n\div a^m=a^{n-m}\]

OpenStudy (anonymous):

Follow these rules?

OpenStudy (unklerhaukus):

yeah for the first expression y^(5/8) ÷ y^(1/2) = y^(5/8 -1/2) =

OpenStudy (unklerhaukus):

for the second expression x^(1/4)*x^1/6*x^(1/3) = x^(1/4 + 1/6 + 1/3) =

OpenStudy (unklerhaukus):

The Index Rules http://mathematicsi.com/rules-of-indices/